OpenStudy (anonymous):

Simplify (x-^2/5y^1/3)15. Write your answer using only positive exponents. 1.) x^6y^5 2.) x^6/y^5 3.)y^5/x^6 4.) 1/x^6y^5

OpenStudy (anonymous):

(x-^2/5y^1/3)15 should this be... (x^-2/5y^1/3)15 or (x^2/5y^1/3)15 ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes what.. which one ?

OpenStudy (anonymous):

(x^-2/5y^1/3)15

OpenStudy (anonymous):

\[\LARGE \left(\frac{x^{-2}}{5y^{\frac13}}\right) \cdot 15=\left(\frac{1}{x^{2}\cdot \sqrt[3]{y}}\right)\cdot \frac{15}{5}=...\]

OpenStudy (anonymous):

so which one is that ?

OpenStudy (anonymous):

you do it...

OpenStudy (anonymous):

...actually it's none of them... check out if you made a mistake typing equations.. or multiple choices !